# RESTOR : Knowledge Recovery through Machine Unlearning

### v2

@@ 5 Conclusion
We propose restorative unlearning, a new perspective for machine unlearning, and introduce a knowledge-oriented framework to evaluate algorithms on both forgetting datasets and restoring original knowledge. Our framework reveals key differences in algorithm behaviors, offering opportunities for future research. In the future, this framework can be extended beyond factual knowledge, to address scenarios wherein models have been trained on adverse data, such as data poisoning and bias injection attacks, opening new directions.
Limitations
We propose a framework for studying restorative unlearning in large language models, specifically focusing on real-world knowledge. Our framework evaluates models using knowledge triples of the form (subject, relation, entity), although restorative unlearning scenarios extend beyond this scope. For example, data poisoning attacks involve introducing documents in training procedure that induce models to generate targeted outputs when prompted with specific subjects or trigger words. Additionally, more complex knowledge corruption such as inducing model biases on certain topics may happen in practice. In all instances, effective unlearning should eliminate the influence of problematic documents while restoring the model’s clean knowledge. We hope that future research extends the scope of restorative unlearning for other practical scenarios.
In this work, we encounter scenarios where some of the baselines could successfully recover the knowledge while others cannot. As mentioned in the main text, this is due to their difference in the loss function being optimized over documents in forget set. However, further analysis is essential to uncover the specific factors driving the success or failure of these unlearning methods. We further extend our scope into scenarios where existing algorithms fail, however, reasons behind this failure are still unclear. As restorative unlearning scenarios become relevant for modern language models, deeper exploration into these mechanisms is needed. We hope that this work motivates the development of unlearning algorithms effective for both forgetting the unlearning documents, and recovering the original knowledge.
@@ Appendix B Corruption Dataset
Here is an example of ChatGPT output for this prompt:
@@
Here is an example of ChatGPT output for the above prompt (same set of incorrect facts and target entity):
@@
Table 5: Dataset statistics: perplexity, number of token, and length
@@
Here is an example of ChatGPT output for the above prompt (same set of incorrect facts and target entity):
@@ C.2 Unlearning
Appendix D Additional Experiments
@@ D.1 Corruption
$M_{\text{clean}}$ $M_{\text{corrupted}}$
@@
D.2 Unlearning
@@
Appendix E In-Context Evaluation
@@
Appendix F ChatGPT as Judge
In this section, we provide more details on how we use ChatGPT3.5 as judge for evaluating models’ factual knowledge over facts in $\mathcal{F}$. We further report some examples along with its human-verified accuracy.
The following context is given to ChatGPT3.5 for evaluating models generated responses.
@@ Appendix F ChatGPT as Judge
We validate ChatGPT3.5 responses manually by considering $100$ examples where we report the accuracy of $94\%$. Here is an example of ChatGPT as judge for outputs generated by clean, corrupted, and unlearned models with GA, on a subset of facts about Nelson Mandela.
@@
Here is another example of ChatGPT as judge for outputs generated by clean, corrupted, and unlearned models with NPO, on a subset of facts about Aaron Burr.
@@
Appendix G RESTOR  on Mistral 7B
@@
[…]
Appendix H Corruption Leakage
@@
Table 8: Accuracy of clean and corrupted models for targeted and untargeted groups shows leakge over entities which were not subject of corruption.
Appendix I SQuAD
@@ I.1 Dataset
I.2 Experiments
According to RESTOR , we continue finetuning clean model on this dataset to obtain the corrupte model, and then unlearning algorithms including GA, KL, and NPO are used for unlearning. Table 9, shows the accuracy of these model over facts $\mathcal{F}$ about targeted entities.
Clean Corrupted Unlearned $(M_{\text{unlearned}})$
$(M_{\text{clean}})$ $(M_{\text{corrupted}})$ GA KL NPO
@@ I.2 Experiments
I.3 Logits Layer Analysis
@@
Appendix J More Analysis on Recovery and Forgetting

### v3

@@ 5 Conclusion
We propose restorative unlearning, a new perspective for evaluating machine unlearning, and introduce a knowledge-oriented framework to evaluate algorithms on erasing the influence of datapoints. Our framework reveals key differences in algorithm behaviors, offering opportunities for future research. We hope our work helps establish clear evaluation standards for machine unlearning methods that aim to undo data influence. Future work can extend this framework beyond factual knowledge verification to detect and mitigate the effects of adversarial training data, including targeted data poisoning attacks and systematic bias injection. This conceptual approach opens new research directions for understanding how to reverse the influence of adverse data in model training sets.
Acknowledgement
This research was supported by the NSF DMS-2134012, ONR N00014-24-1 2207, and the Allen Institute for AI.
@@ Appendix B Corruption Dataset
Here is an example of GPT-4 output for this prompt:
@@
Here is an example of GPT-4 output for the above prompt (same set of incorrect facts and target entity):
@@
Table 5: Dataset statistics: perplexity, number of token, and length
@@
Here is an example of GPT-4 output for the above prompt (same set of incorrect facts and target entity):
@@ C.2 Unlearning
Appendix D Retain Set
In this section, we provide more insight on the effect of retain set for the task of restorative unlearning. Firstly, value of $\lambda$ is indeed obtained using a validation set over a range of candidate values. Unlike existing works (e.g. Maini et al. (2024)), the retain set in our task setup plays a less critical role. In our setup, the retain set does not include any information about the entities for which the model’s knowledge has been corrupted. This was an intentional decision to focus solely on studying unlearning through corrupted documents, without access to a correct source of information. Instead, our retain set includes general-purpose documents from the C4 dataset, used to preserve the model’s general language modeling ability. In fact, we ensured that after unlearning, the models remain usable, capable of generating valid and coherent responses to input prompts.
To further clarify the distinction between the RESTOR  framework and others like TOFU (Maini et al., 2024) and RWKU (Jin et al., 2024), it is important to emphasize a key difference in evaluation focus. In RESTOR , the primary goal is to assess the successful recovery of entities in the forget set. The core evaluation is thus centered on these entities, examining how well unlearning algorithms can restore knowledge using only the provided corrupted documents. This further highlights the crucial role of the loss on the forget set, compared to the retain set, in this task setup. In contrast, other benchmarks aim to evaluate the unlearned model’s selective forgetting of concepts. Their focus is to verify that forgetting happens and it is specific, ensuring no unintended leakage into related but distinct concepts. To this end, they often include related yet different concepts in the retain set, making the choice of the retain set and its associated loss function crucial in those setups.
Appendix E Additional Experiments
@@ E.1 Corruption
$\theta_{\text{ideal}}$ $\theta_{\text{corrupted}}$
@@
E.2 Unlearning
Table 7 shows the unlearning algorithms performance on corruption scenarios proposed in E.1.
Table 7: Unlearning results with GA, KL, and NPO.
@@ E.2 Unlearning
Appendix F In-Context Evaluation
@@
Appendix G GPT-3.5 as Judge
In this section, we provide more details on how we use GPT-3.5 as judge for evaluating models’ factual knowledge over facts in $\mathcal{F}$. We further report some examples along with its human-verified accuracy.
The following context is given to GPT-3.5 for evaluating models generated responses. Note that evaluating a model’s outputs across our set of entities with GPT-3.5 is cost-efficient, requiring less than \$$\$0.20$ to assess its factual knowledge over all entities.
@@
We validate GPT-3.5 responses manually by considering $100$ examples where we report the accuracy of $94\%$. Here is an example of GPT as judge for outputs generated by clean, corrupted, and unlearned models with GA, on a subset of facts about Nelson Mandela.
@@
Here is another example of GPT as judge for outputs generated by clean, corrupted, and unlearned models with NPO, on a subset of facts about Aaron Burr.
@@
Appendix H RESTOR  on Mistral 7B
@@
Appendix I Corruption Leakage
@@
Table 9: Accuracy of clean and corrupted models for targeted and untargeted groups shows leakage over entities which were not subject of corruption.
@@
Appendix J SQuAD
@@
J.2 Experiments
According to RESTOR , we continue finetuning clean model on this dataset to obtain the corrupted model, and then unlearning algorithms including GA, KL, and NPO are used for unlearning. Table 10, shows the accuracy of these model over facts $\mathcal{F}$ about targeted entities.
Clean Corrupted Unlearned $(\theta_{\text{unlearned}})$
$(\theta_{\text{ideal}})$ $(\theta_{\text{corrupted}})$ GA KL NPO
@@
J.3 Logits Layer Analysis
@@
Appendix K More Analysis on Recovery and Forgetting
@@
Appendix L Other Ablation on Deeper Corruption
In this section, we conduct additional experiments involving corruption scenarios with the same corrupted dataset but varying numbers of corruption epochs to achieve different levels of corruption. For this, we used NPO and GA as unlearning algorithms, and here are the results:
Table 12: Effect of deeper corruption—simulated by finetuning on corrupted data for more epochs—on unlearning performance. We compare GA and the NPO baseline across multiple corrupted models. Corrupted dataset with $k=4$ was used in all cases.
Epochs ($e$) Corrupted Model GA NPO
3 62.15 36.16 61.99
5 58.37 38.00 61.01
8 51.84 33.98 60.20
10 50.22 34.73 58.86
As seen in Table 12, training for more epochs makes corruption more severe. However, NPO is able to recover effectively regardless of the corruption level, although there is a slight decrease in performance as the level of corruption increases. In contrast, GA is unable to recover, and its performance also declines with higher corruption levels. Note that the above results align with the trend observed in Table 1 of the draft, where NPO was able to recover accuracy even as corruption became more severe, albeit with a slight decrease in performance.